# Learning Representations for Independence Testing

### v2

@@ 4.1 HSIC as a lower bound on MI
Suppose the kernel $h$ on $(x,y)$ pairs of Proposition 3.1 is bounded: $\sup_{x\in\mathcal{X},y\in\mathcal{Y}}h((x,y),(x,y))\leq\nu^{2}$ for $\nu\geq 0$. Then
@@
by the reproducing property and Cauchy-Schwarz, so $\lVert f\rVert_{\mathcal{H}}\geq\frac{1}{\nu}\sup_{x,y}\lvert f(x,y)\rvert=\frac{1}{\nu}\lVert f\rVert_{\infty}$. This implies that $\{f:\lVert f\rVert_{\mathcal{H}}\leq 1\}\subseteq\{f:\lVert f\rVert_{\infty}\leq\nu\}$, and so by Proposition 3.1,
@@
In the setting of Proposition 3.1, suppose $\sup_{x\in\mathcal{X},y\in\mathcal{Y}}h((x,y),(x,y))\leq\nu^{2}$. Then
@@
The second bound is tighter for large values of $\operatorname{I}(X;Y)$, but both are monotonic in $\operatorname{HSIC}(X,Y)/\nu^{2}$. We could thus consider, as in Section 2, choosing kernels $k$, $l$ to maximize a lower bound on $\operatorname{I}(X;Y)$, by maximizing $\operatorname{HSIC}(X,Y)/\nu^{2}$. Indeed, maximizing HSIC has been used by previous applications in many areas (e.g. Blaschko & Gretton, 2009; Song et al., 2012; Li et al., 2021; Dong et al., 2023).
@@ 4.2 Kernel-based tests and variational MI tests
Now consider a test based on $\operatorname{MMD}_{f}(\mathbb{P}_{xy},\mathbb{P}_{x}\times\mathbb{P}_{y})$ using a kernel of the form $h((x,y),(x^{\prime},y^{\prime}))=f(x,y)f(x^{\prime},y^{\prime})$ for some real-valued function $f$. If $f(x,y)=f_{1}(x)f_{2}(y)$, this is an HSIC test with kernels $k(x,x^{\prime})=f_{1}(x)f_{1}(x^{\prime})$ and $l(y,y^{\prime})=f_{2}(y)f_{2}(y^{\prime})$. Because $\phi_{h}(x,y)=f(x,y)\in\mathbb{R}$ is a valid feature map, every function in $\mathcal{H}$ is of the form $\alpha f=[(x,y)\mapsto\alpha f(x,y)]$ with $\lVert\alpha f\rVert_{\mathcal{H}}=\lvert\alpha\rvert$. By Proposition 3.1,
$\operatorname{MMD}_{f}(\mathbb{P}_{xy},\mathbb{P}_{x}\times\mathbb{P}_{y})^{2}=\left(\sup_{\lvert\alpha\rvert\leq 1}\alpha\left(\operatorname*{\mathbb{E}}f(X,Y)-\operatorname*{\mathbb{E}}f(X,Y^{\prime})\right)\right)^{2}=\bigl{(}\operatorname*{\mathbb{E}}f(X,Y)-\operatorname*{\mathbb{E}}f(X,Y^{\prime})\bigr{)}^{2}.$
The plug-in estimator corresponding to the biased HSIC estimator (4) would yield the test statistic
$\widehat{\operatorname{MMD}}^{2}_{\rm b}(\mathbb{X},\mathbb{Y})=\left(\frac{1}{m}\sum_{i=1}^{m}f(x_{i},y_{i})-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}f(x_{i},y_{j})\right)^{2}.$ (8)
If $f(x,y)=f_{1}(x)f_{2}(y)$, this is exactly $\widehat{\operatorname{HSIC}}_{\rm b}$.
@@ 5 Learning Representations for Independence Testing
$\Pr_{\mathfrak{H}_{1}}(\hat{T}>r_{m})\sim\Phi\left(\frac{\sqrt{m}(T-T_{\mathfrak{H}_{0}})}{\tau_{\mathfrak{H}_{1}}}-\frac{\tau_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{1}}}\Phi^{-1}(1-\alpha)\right)\;\text{and}\;\Pr_{\mathfrak{H}_{1}}(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m})\sim\Phi\left(\frac{\sqrt{m}\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}-\frac{\Psi^{-1}(1-\alpha)}{\sqrt{m}\sigma_{\mathfrak{H}_{1}}}\right).$
In both cases, as the sample size $m$ grows, the power is dominated by the signal-to-noise ratio which we define as
@@
Maximizing the SNR maximizes the limiting power of the test as $m\to\infty$. Thus, maximizing estimates of the SNR (based on a finite number of samples $n$) will roughly maximize the power of a test at an arbitrarily large sample size ($m\to\infty$); also see discussion by Sutherland & Deka (2019); Deka & Sutherland (2023).
Alternatively, we can also choose to include the threshold-dependent term. This may be helpful, particularly in estimating the power at points where the number of samples is not overwhelming: for tests with a 50% probability of rejection, the two terms in the power expression are necessarily the same size. For NDS, the gap between the two terms only grows with $\sqrt{m}$, and additionally estimating $\tau_{\mathfrak{H}_{1}}$ is no problem when already estimating $T_{\mathfrak{H}_{1}}$, thus we propose to include the threshold term in choosing an NDS critic. For HSIC, however, the threshold-dependent term is both relatively less important as $m$ grows and much more difficult to estimate. [^5] Ren et al. (2024) proposed using a moment-matched gamma approximation to the threshold, and claimed that not including this threshold can lead to catastrophically wrong kernel choice. We argue in Section E.5, however, that their argument is unjustified. We also find experimentally in Section 6 that including it can choose worse kernels in practice. More common contemporary approaches to estimating the threshold use permutation testing or eigendecomposition; both involve substantial computational overhead, and the latter is particularly expensive to differentiate while the former can be quite difficult to usefully estimate on the same data as the HSIC estimate, as observed for MMD by Deka & Sutherland (2023) and discussed further in Section 7. Therefore, we propose to optimize the SNR with the threshold term for NDS, and without it for HSIC. Our population level objectives, as functions of random variables $X$ and $Y$, are then
@@
which is just a regularized version of the variance estimate stated in Proposition A.1. Following Song et al. (2012), this can be computed more efficiently with $R=\frac{((n-4)!)^{2}}{4n((n-1)!)^{2}}\lVert\boldsymbol{h}\rVert^{2}$, where the vector $\boldsymbol{h}$ is
@@
Maximizing (12) or (13) then gives us a structured approach to select the critic or kernel yielding the strongest asymptotic test. Typically, we run some variant of a gradient-based optimization algorithm with respect to the parameters of the critic or kernel class.
@@
     $\hat{J}^{\lambda}(\mathbb{X},\mathbb{Y};\theta)\leftarrow$ compute SNR estimate; # as in Equation 12 or Equation 13
     $\theta\leftarrow\theta+\eta\nabla_{\theta}\hat{J}^{\lambda}(\mathbb{X},\mathbb{Y};\theta)$ gradient ascent step
@@
[…]
@@
The overall procedure, written based on minibatch gradient ascent for simplicity, is shown in Algorithm 1. In practice, we use AdamW (Loshchilov & Hutter, 2019), and draw minibatches in epochs; experimental details are given in Section D.1. The randomized $p$-value is potentially conservative in the case of ties; breaking ties randomly can give a slight improvement when ties are present (Hemerik & Goeman, 2018).
@@
While neither $\hat{J}_{\text{NDS}}^{\lambda}$ nor $\hat{J}_{\text{HSIC}}^{\lambda}$ are biased estimates of the relevant population quantities, [^8] Indeed, no unbiased estimator is likely to exist; see Appendix A of Deka & Sutherland 2023. the following Theorems 5.1 and 5.2 show that they are both uniformly bounded in probability.
@@
Let $\{f_{\theta}:\theta\in\Theta\}$ be a critic family with parameter space $\Theta$ satisfying Assumptions (A), (B) and (C) in Section B.2 which define a critic bound $B$, dimension $D$, and smoothness $L$. Suppose $\tau_{\theta}^{2}\geq s^{2}$ for some positive $s$ under both $\mathfrak{H}_{0}$ and $\mathfrak{H}_{1}$, and $\lambda=\Theta(n^{-1/3})$. Then
@@
Let $\{k_{\omega}:\omega\in\Omega\}$ and $\{l_{\gamma}:\gamma\in\Gamma\}$ be kernel families with parameter spaces $\Omega$ and $\Gamma$ satisfying Assumptions (A’), (B’) and (C’) in Section C.2 which define kernel dimensions $D_{\Omega}$, $D_{\Gamma}$ and smoothnesses $L_{k}$, $L_{l}$. Suppose $\sigma_{\omega,\gamma}^{2}\geq s^{2}>0$ under $\mathfrak{H}_{1}$ and $\lambda=\Theta(n^{-1/3})$. Then
@@
Appendices B and C state and prove non-asymptotic versions of the above results; the assumptions and proof techniques are based on those of Liu et al. (2020). [^9] Ren et al. (2024) stated a similar result to Theorem 5.2, but with an incorrect proof; although they claim uniform convergence over their threshold estimate, their proof makes no attempt at showing uniformity, instead showing pointwise convergence.
Following Theorem B.13 (a standard result from van der Vaart, 1998), successfully maximizing these estimates will thus also maximize the population quantity, and consequently the asymptotic test power, for sufficiently large training set sizes $n$.
@@ 6 Experiments
  • HSIC-D: HSIC using deep kernels on each space $\mathcal{X}$ and $\mathcal{Y}$; simultaneously trained via Section 5.
  • HSIC-Dx: HSIC using a tied deep kernel, i.e. $k_{\omega}=l_{\gamma}$, and trained via Section 5.
  • HSIC-O: HSIC using Gaussian kernels, with each bandwidth parameter optimized via Section 5.
  • NDS: The neural dependency statistic (2) trained via Section 5.
@@
$[X_{1},Y_{\left\lfloor{d/2}\right\rfloor},\dots,X_{\left\lceil{d/2}\right\rceil},Y_{1}]\sim\sum_{i=1}^{2}\frac{1}{2}\mathcal{N}\left(\textbf{0}_{d},\begin{bmatrix}1&0.5(-1)^{i}&\textbf{0}^{T}_{d-2}\\ 0.5(-1)^{i}&1&\textbf{0}^{T}_{d-2}\\ \textbf{0}_{d-2}&\textbf{0}_{d-2}&I_{d-2}\end{bmatrix}\right),$
@@
We first compare how well methods identify dependency with a large-size training set, by comparing the rate at which the learned tests achieve perfect power (1.0) as the test size $m$ increases. Training and validation sizes for each dataset are given in Section D.1, and results are visualized in Figure 4 (with comprehensive results in Figure 9). Overall, HSIC-D outperforms baselines, and is able to reach perfect power at smaller test sizes $m$. We note that HSIC-O performs reasonably well for simpler problems like Sinusoid, but struggles on harder problems like RatInABox. This suggests that no Gaussian kernel on the input space is well-suited for the task. On the other hand, kernels applied to optimal representations of the data (HSIC-D) are the most powerful. Surprisingly, optimizing the power of NDS tests is significantly weaker than directly maximizing InfoNCE or NWJ; we elaborate on this in the following section.
@@
These are baselines considered in this paper. See more details in Section 6.
@@
A drawback to kernel selection via optimization is that we must hold out a split of the data for training. In contrast, HSIC-M and HSIC-Agg are able to utilize the entirety of the dataset for their test. To examine this trade-off, we consider consistent data splitting at datasets of varying sizes. For the HDGM and Sinusoid problems we use a 7:2:1 train-val-test split, and for RatInABox and Wine we use a 3:2 train-test split. Conversely, HSIC-M and HSIC-Agg use the entire dataset for testing. Results are shown in Figure 5. Non-splitting methods are able to take advantage of the additional test samples and outperform some data-splitting methods at smaller dataset sizes; for large dataset sizes the reverse is true.
@@
These are baselines considered in this paper. See more details in Section 6.
@@
Our HSIC objective $\hat{J}_{\operatorname{HSIC}}^{\lambda}$ disregards the threshold-dependent term, whereas the NDS objective $\hat{J}_{\text{NDS}}^{\lambda}$ does not. The impact this term has during training, however, is still unclear. We compare the strength of tests maximized with and without this term in Figure 15. It is important to note that the threshold estimate for HSIC is challenging (see also Footnote 5); we use a method based on the moment-matched Gamma approximation, but with improved gradient estimation rather than the finite differences estimate they used (Section E.6). This estimate, in either form, introduces substantial additional code complexity and its own bias and variance that could each hurt overall performance. In general, we find that including the threshold for HSIC has no strong effect on simpler problems, but degrades the test power for harder problems. On the other hand, NDS with the threshold yields stronger tests. This is not too surprising as, unlike for HSIC, the null distribution here follows the CLT and is easily approximated given enough samples.

### v3

@@ 4.1 HSIC as a lower bound on MI
Suppose the kernel $h$ on $(x,y)$ pairs of Proposition˜3.1 is bounded: $\sup_{x\in\mathcal{X},y\in\mathcal{Y}}h((x,y),(x,y))\leq\nu^{2}$ for $\nu\geq 0$. Then
@@
by the reproducing property and Cauchy-Schwarz, so $\lVert f\rVert_{\mathcal{H}}\geq\frac{1}{\nu}\sup_{x,y}\lvert f(x,y)\rvert=\frac{1}{\nu}\lVert f\rVert_{\infty}$. This implies that $\{f:\lVert f\rVert_{\mathcal{H}}\leq 1\}\subseteq\{f:\lVert f\rVert_{\infty}\leq\nu\}$, and so by Proposition˜3.1,
@@
In the setting of Proposition˜3.1, suppose $\sup_{x\in\mathcal{X},y\in\mathcal{Y}}h((x,y),(x,y))\leq\nu^{2}$. Then
@@
The second bound is tighter for large values of $\operatorname{I}(X;Y)$, but both are monotonic in $\operatorname{HSIC}(X,Y)/\nu^{2}$. We could thus consider, as in Section˜2, choosing kernels $k$, $l$ to maximize a lower bound on $\operatorname{I}(X;Y)$, by maximizing $\operatorname{HSIC}(X,Y)/\nu^{2}$. Indeed, maximizing HSIC has been used by previous applications in many areas (e.g. Blaschko & Gretton, 2009; Song et al., 2012; Li et al., 2021; Dong et al., 2023).
@@ 4.2 Kernel-based tests and variational MI tests
Now consider a test based on $\operatorname{MMD}_{f}(\mathbb{P}_{xy},\mathbb{P}_{x}\times\mathbb{P}_{y})$ using a kernel of the form $h((x,y),(x^{\prime},y^{\prime}))=f(x,y)f(x^{\prime},y^{\prime})$ for some real-valued function $f$. If $f(x,y)=f_{1}(x)f_{2}(y)$, this is an HSIC test with kernels $k(x,x^{\prime})=f_{1}(x)f_{1}(x^{\prime})$ and $l(y,y^{\prime})=f_{2}(y)f_{2}(y^{\prime})$. Because $\phi_{h}(x,y)=f(x,y)\in\mathbb{R}$ is a valid feature map, every function in $\mathcal{H}$ is of the form $\alpha f=[(x,y)\mapsto\alpha f(x,y)]$ with $\lVert\alpha f\rVert_{\mathcal{H}}=\lvert\alpha\rvert$. By Proposition˜3.1,
$\operatorname{MMD}_{f}(\mathbb{P}_{xy},\mathbb{P}_{x}\times\mathbb{P}_{y})^{2}=\left(\sup_{\lvert\alpha\rvert\leq 1}\alpha\left(\operatorname*{\mathbb{E}}_{\mathbb{P}_{xy}}f(X,Y)-\operatorname*{\mathbb{E}}_{\mathbb{P}_{x}\times\mathbb{P}_{y}}f(X,Y^{\prime})\right)\right)^{2}=\bigl(\operatorname*{\mathbb{E}}_{\mathbb{P}_{xy}}f(X,Y)-\operatorname*{\mathbb{E}}_{\mathbb{P}_{x}\times\mathbb{P}_{y}}f(X,Y^{\prime})\bigr)^{2}.$
The plug-in estimator would yield the test statistic
$\widehat{\operatorname{MMD}}^{2}_{\rm b}(\mathbb{X},\mathbb{Y})=\left(\frac{1}{m}\sum_{i=1}^{m}f(x_{i},y_{i})-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}f(x_{i},y_{j})\right)^{2},$ (8)
which when $f(x,y)=f_{1}(x)f_{2}(y)$, corresponds exactly to $\widehat{\operatorname{HSIC}}_{\rm b}$.
@@ 5 Learning Representations for Independence Testing
$\Pr_{\mathfrak{H}_{1}}(\hat{T}>r_{m})\approx\Phi\left(\frac{\sqrt{m}(T-T_{\mathfrak{H}_{0}})}{\tau_{\mathfrak{H}_{1}}}-\frac{\tau_{\mathfrak{H}_{0}}}{\tau_{\mathfrak{H}_{1}}}\Phi^{-1}(1-\alpha)\right)\;\text{and}\;\Pr_{\mathfrak{H}_{1}}(\widehat{\operatorname{HSIC}}_{\rm u}>r_{m})\approx\Phi\left(\frac{\sqrt{m}\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}-\frac{\Psi^{-1}(1-\alpha)}{\sqrt{m}\sigma_{\mathfrak{H}_{1}}}\right).$
In both cases, as the sample size $m$ grows, the power is dominated by the signal-to-noise ratio
@@
Maximizing the SNR maximizes the limiting power of the test as $m\to\infty$. (Also see Appendix˜B.) Thus, maximizing estimates of the SNR (based on a finite number of samples $n$) will roughly maximize the power of a test at an arbitrarily large sample size ($m\to\infty$); also see discussion by Sutherland & Deka (2019); Deka & Sutherland (2023).
Alternatively, we can also choose to include the threshold-dependent term. This may be helpful, particularly in estimating the power at points where the number of samples is not overwhelming: for tests with a 50% probability of rejection, the two terms in the power expression are necessarily the same size. For NDS, the gap between the two terms only grows with $\sqrt{m}$, and additionally estimating $\tau_{\mathfrak{H}_{1}}$ is no problem when already estimating $T_{\mathfrak{H}_{1}}$, thus we propose to include the threshold term in choosing an NDS critic. For HSIC, however, the threshold-dependent term is both relatively less important as $m$ grows and much more difficult to estimate. [^6] Ren et al. (2024) proposed using a moment-matched gamma approximation to the threshold, and claimed that not including this threshold can lead to catastrophically wrong kernel choice. We argue in Section F.5, however, that their argument is unjustified. We also find experimentally in Section 6 that including it can choose worse kernels in practice. More common contemporary approaches to estimating the threshold use permutation testing or eigendecomposition; both involve substantial computational overhead, and the latter is particularly expensive to differentiate while the former can be quite difficult to usefully estimate on the same data as the HSIC estimate, as observed for MMD by Deka & Sutherland (2023) and discussed further in Section 7. Therefore, we propose to optimize the SNR with the threshold term for NDS, and without it for HSIC. Our population level objectives, as functions of random variables $X$ and $Y$, are then
@@
which is just a regularized version of the variance estimate stated in Proposition˜A.1. Following Song et al. (2012), this can be computed more efficiently with $R=\frac{((n-4)!)^{2}}{4n((n-1)!)^{2}}\lVert\boldsymbol{h}\rVert^{2}$, where the vector $\boldsymbol{h}$ is
@@
Maximizing (12) or (13) then gives us a structured approach to select the critic or kernel yielding the strongest asymptotic test. Typically, we run some variant of a gradient-based optimization algorithm with respect to the parameters of the critic or kernel class, and do hyperparameter selection and early stopping based on the same objective on a validation set.
@@
   $\hat{J}^{\lambda}(\mathbb{X},\mathbb{Y};\theta)\leftarrow$ compute SNR estimate; # as in Equation˜12 or Equation˜13
   $\theta\leftarrow\theta+\eta\nabla_{\theta}\hat{J}^{\lambda}(\mathbb{X},\mathbb{Y};\theta)$ # gradient ascent step
@@
  $\mathit{perm}_{1}\leftarrow{\mathcal{T}}(S^{\rm te}_{X},S^{\rm te}_{Y};\theta)$ # evaluate test statistic
  for $i=2,\dots,n_{\mathit{perm}}$ do
   $\mathit{perm}_{i}\leftarrow\mathcal{T}(S^{\rm te}_{X},\operatorname{shuffle}(S^{\rm te}_{Y});\theta)$ # no need to shuffle both sets
  Output: $k_{\omega}$, $l_{\gamma}$, $\mathit{perm}_{1}$, $p$-value $\frac{1}{n_{\mathit{perm}}}\sum_{i=1}^{n_{\mathit{perm}}}\mathbbm{1}(\mathit{perm}_{i}\geq\mathit{perm}_{1})$
@@
The overall procedure, written based on minibatch gradient ascent for simplicity, is shown in Algorithm˜1. In practice, we use AdamW (Loshchilov & Hutter, 2019), and draw minibatches in epochs; experimental details are given in Section˜E.1. The randomized $p$-value is potentially conservative in the case of ties; breaking ties randomly can give a slight improvement when ties are present (Hemerik & Goeman, 2018).
@@
While neither $\hat{J}_{\text{NDS}}^{\lambda}$ nor $\hat{J}_{\text{HSIC}}^{\lambda}$ are biased estimates of the relevant population quantities, [^9] Indeed, no unbiased estimator is likely to exist; see Appendix A of Deka & Sutherland 2023. the following Theorems˜5.1 and 5.2 show that they are both uniformly bounded in probability.
@@
Let $\{f_{\theta}:\theta\in\Theta\}$ be a critic family with parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C) in Section˜C.2 which define a critic bound $B$, dimension $D$, and smoothness $L$. Suppose $\tau_{\theta}^{2}\geq s^{2}$ for some positive $s$ under both $\mathfrak{H}_{0}$ and $\mathfrak{H}_{1}$, and $\lambda=\Theta(n^{-1/3})$. Then
@@
Let $\{k_{\omega}:\omega\in\Omega\}$ and $\{l_{\gamma}:\gamma\in\Gamma\}$ be kernel families with parameter spaces $\Omega$ and $\Gamma$ satisfying Assumptions˜(A’), (B’) and (C’) in Section˜D.2 which define kernel dimensions $D_{\Omega}$, $D_{\Gamma}$ and smoothnesses $L_{k}$, $L_{l}$. Suppose $\sigma_{\omega,\gamma}^{2}\geq s^{2}>0$ under $\mathfrak{H}_{1}$ and $\lambda=\Theta(n^{-1/3})$. Then
@@
Appendices˜C and D state and prove non-asymptotic versions of the above results based on covering numbers; the assumptions and proof techniques are similar to those of Liu et al. (2020) [^10] Ren et al. (2024) stated a similar result to Theorem 5.2, but with a seemingly incorrect proof: in their Appendix D.3, the term $\lvert r^{(n)}-r\rvert$ in their (63) should be bounded uniformly over kernel parameters, but they appeal to a result about convergence of distribution functions for a single distribution. For the uniform result over kernel parameters that their theorem statement claims, they would need to additionally use a cover or other approach (as is done for the HSIC and the variance estimators), but none is attempted., but is slightly more involved for a few reasons: 1) Uniform convergence of HSIC and its variance estimators now include two kernels, $k$ and $l$, instead of one, and a somewhat different estimator form. 2) For NDS we also need to show convergence of a null-to-alternative variance ratio, a term which is not present in the MMD or HSIC objectives.
Following Theorem˜C.13 (a standard result from van der Vaart, 1998), successfully maximizing these estimates will thus also maximize the population quantity, and consequently the asymptotic test power, for sufficiently large training set sizes $n$.
@@ 6 Experiments
The repository https://github.com/xunathan96/deephsic/ contains implementations of our methods and code for our experiments.
@@
  • HSIC-D: HSIC using deep kernels on each space $\mathcal{X}$ and $\mathcal{Y}$; simultaneously trained via Section˜5.
  • HSIC-Dx: HSIC using a tied deep kernel, i.e. $k_{\omega}=l_{\gamma}$, and trained via Section˜5.
  • HSIC-O: HSIC using Gaussian kernels, with each bandwidth parameter optimized via Section˜5.
  • NDS: The neural dependency statistic (2) trained via Section˜5.
@@
[…]
@@
We first compare how well methods identify dependency with a large-size training set, by comparing the rate at which the learned tests achieve perfect power (1.0) as the test size $m$ increases. Training and validation sizes for each dataset are given in Section˜E.1, and results are visualized in Figure˜4 (with comprehensive results in Figure˜9). Overall, HSIC-D outperforms baselines, and is able to reach perfect power at smaller test sizes $m$. We note that HSIC-O performs reasonably well for simpler problems like Sinusoid, but struggles on harder problems like RatInABox. This suggests that no Gaussian kernel on the input space is well-suited for the task. On the other hand, kernels applied to optimal representations of the data (HSIC-D) are the most powerful. Surprisingly, optimizing the power of NDS tests is significantly weaker than directly maximizing InfoNCE or NWJ; we elaborate on this in the following section.
@@
[…]
@@
A drawback to kernel selection via optimization is that we must hold out a split of the data for training. In contrast, HSIC-M and HSIC-Agg are able to utilize the entirety of the dataset for their test. To examine this trade-off, we consider consistent data splitting at datasets of varying sizes. For the HDGM and Sinusoid problems we use a 7:2:1 train-val-test split, and for RatInABox and Wine we use a 3:2 train-test split. Conversely, HSIC-M and HSIC-Agg use the entire dataset for testing. Results are shown in Figure˜5. Non-splitting methods are able to take advantage of the additional test samples and outperform some data-splitting methods at smaller dataset sizes; for large dataset sizes the reverse is true. The validation set in this experiment is used only for early stopping of the training process, and not for hyperparameter selection.
@@
[…]
@@
Our HSIC objective $\hat{J}_{\operatorname{HSIC}}^{\lambda}$ disregards the threshold-dependent term, whereas the NDS objective $\hat{J}_{\text{NDS}}^{\lambda}$ does not. The impact this term has during training, however, is still unclear. We compare the strength of tests maximized with and without this term in Figure˜15. It is important to note that the threshold estimate for HSIC is challenging (see also Footnote˜6); we use a method based on the moment-matched Gamma approximation, but with improved gradient estimation rather than the finite differences estimate they used (Section˜F.6). This estimate, in either form, introduces substantial additional code complexity and its own bias and variance that could each hurt overall performance. In general, we find that including the threshold for HSIC has no strong effect on simpler problems, but degrades the test power for harder problems. On the other hand, NDS with the threshold yields stronger tests. This is not too surprising as, unlike for HSIC, the null distribution here follows the CLT and is easily approximated given enough samples.